OpenStudy (anonymous):

Determine whether the lines through the pairs of points are parallel, perpendicular, or neither. 6x + 2y = 8 15x + 5y = 21

OpenStudy (amistre64):

6(5) - 15(2) = 0, so they are parallel by some standard

OpenStudy (amistre64):

-6/2 = -3 -15/5 = -3 same slopes